# Network reconstruction via the minimum description length principle

### v1

@@ V.3 Node parameters
$P(\bm{\theta}|\lambda_{\theta},\Delta_{\theta})\\ =\frac{\prod_{k}n_{k}!\times\mathrm{e}^{-\lambda\sum_{k}|u_{k}|}\sinh(\lambda_{\theta}\Delta_{\theta})^{K_{\theta}-\mathds{1}_{0\in\bm{u}}}(1-\mathrm{e}^{-\lambda_{\theta}\Delta_{\theta}})^{\mathds{1}_{0\in\bm{u}}}}{N!{N-1\choose K_{\theta}-1}N},$ (34)
@@ V.4 Empirical evaluation, contrast to $L_{1}$ regularization and decimation
In Fig. 3 we show the result of the reconstruction using both our MDL scheme and $L_{1}$ regularization together with 5-fold cross-validation. Similarly to the Karate Club example of Fig. 1, the $L_{1}$ inference incorporates one order of magnitude more non-zero edges ($E=10,911$) when compared with the MDL method ($E=1,540$). The edges obtained with $L_{1}$ are mostly of low magnitude (see Fig. 3c), and based on our analysis for synthetic data, are much more likely to be overfitting. We can find evidence of this by performing cross-validation also with the MDL method, as shown in Fig. 3d, which yields a significantly improved predictive performance of the held-out data. Both the predictive performance and description length improve when performing MDL regularization together with community detection, in accordance with Ref. [54].
@@
The interpretation of the two inferred models is fairly different. Although they share similar large-scale structures, the one inferred by $L_{1}$ possesses weights that are often negative, indicating some amount of direct antagonism between deputies, i.e. conditioning on the “yes” vote of one deputy will increase the probability of a “no” vote of another. Instead, the model inferred by MDL possesses only positive weights, divided in three categories, indicating an overwhelming tendency towards selective agreement, instead of direct antagonism. The groups it finds also align very well with the available metadata: the largest group corresponds to the governing coalition, whereas the remaining are opposition parties forming fragmented blocks.
@@
Figure 4: Decimation procedure [29] employed on the same data as Fig. 3. The bottom panel shows the maximum likelihood growing monotonically as a function of the number of nonzero edges considered during decimation, and the top shows the similarity (weighted and binarized) with the network inferred via MDL.
@@
We take the opportunity to compare with another regularization scheme called “decimation”, proposed in Ref. [29]. That approach is not based on shrinkage, and instead proceeds by first performing an unregularized maximum likelihood, resulting in a full network with no zero weights, then forcing a small number of the weights with smallest magnitude to zero, and proceeding recursively on the remaining entries in the same manner, until sufficiently many entries have been set to zero. Besides the increased computational cost of inferring the weights multiple times, and starting with a full network—thus imposing an overall complexity of at least $O(N^{2})$—this approach offers no obvious, principled criterion to determine when to stop the decimation. In Ref. [29] a heuristic stopping criterion was proposed by observing a point at which the likelihood no longer changes significantly in synthetic examples sampled from the true model. However, there is no guarantee that such a signal will exist in empirical data. In Fig. 4 we show for our Brazilian congress example the model likelihood offers no clear indication of when to stop the decimation and how many edges should be inferred in the end. The Jaccard similarity with the network inferred via MDL remains low through the whole procedure, peaking at close to $1/2$ for the binarized version when both networks have similar number of edges. Due to these limitations, we see no advantage of using decimation over our approach based on MDL.
@@ VI Empirical case study: network of microbial interactions
Figure 5: Reconstructed networks for the (a) human microbiome project (HMP) with $N=45,383$, $M=4,788$, and $E=122,804$ and (b) earth microbiome project (EMP) with $N=126,730$, $M=23,323$, and $E=735,868$, using the MDL regularization method. The top panels show the networks obtained, with edge weights indicated as colors. The middle panel shows the OTU counts in (a) different body sites and (b) earth habitats. The bottom panel shows the edge weight distributions (with the same colors as the top panel), the node field distributions, the degree distributions (where node $i$ has total degree $k_{i}=\sum_{j}A_{ij}$, as well as positive and negative degrees, $k^{+}=\sum_{j}A_{ij}\mathds{1}_{W_{ij}>0}$ and $k^{-}=\sum_{j}A_{ij}\mathds{1}_{W_{ij}<0}$, respectively), the node strength distributions (where node $i$ has total strength $d_{i}=\sum_{j}W_{ij}$, as well as positive and negative strenght, $d_{i}^{+}=\sum_{j}W_{ij}\mathds{1}_{W_{ij}>0}$ and $d_{i}^{-}=\sum_{j}W_{ij}\mathds{1}_{W_{ij}<0}$, respectively), and the average node fields as a function of the degrees (with the expected “magnetization” in the inset).
@@
using the MDL regularization for both $\bm{W}$ and $\bm{\theta}$ discussed previously. (We will not include a comparison with $L_{1}$, since the cross-validation procedure it requires becomes computationally prohibitive for problems of this magnitude.)
@@ VI.0.1 Predicting stability and outcome of interventions
In Fig. 7 we show the distribution of $z(j,\alpha)$ for macrostates simulated with the Ising model inferred from the HMP data, using the belief-propagation (BP) method [65]. Although most perturbations are inconsequential, we observe that sometimes hundreds of species can disappear (or even appear) after a single species is removed.
@@ Appendix B Species taxonomies
Figure 9: Taxonomic classification of species for the human microbiome project (HMP), overlaid on the inferred networks of Fig. 5, together with the total species count over all samples.
@@
Figure 10: Taxonomic classification of species for the earth microbiome project (EMP), overlaid on the inferred networks of Fig. 5, together with the total species count over all samples.
In Figs. 9 and 10 we show the taxonomic classification of the species for the reconstructed networks of Fig. 5. In both cases we observe that most groups inferred with the SBM incorporated in our regularization scheme contain one dominating taxonomic category, at all levels, although a single category is often split into many groups.

### v3

@@ V.2 Joint reconstruction with community detection
We refer the reader to Ref. [70] for more details about the SBM formulation, and to Ref. [68] for its effect in network reconstruction.
@@ V.3 Node parameters
$P(\bm{\theta}|\lambda_{\theta},\Delta_{\theta})=\prod_{k}n_{k}!\times\mathrm{e}^{-\lambda\sum_{k}|u_{k}|}\\ \times\frac{\sinh(\lambda_{\theta}\Delta_{\theta})^{K_{\theta}-\mathds{1}_{0\in\bm{u}}}(1-\mathrm{e}^{-\lambda_{\theta}\Delta_{\theta}})^{\mathds{1}_{0\in\bm{u}}}}{N!{N-1\choose K_{\theta}-1}N},$ (34)
@@ V.4 Empirical evaluation, contrast to $L_{1}$ regularization and decimation
In Fig. 3 we show the result of the reconstruction using both our MDL scheme and $L_{1}$ regularization together with 5-fold cross-validation. Similarly to the Karate Club example of Fig. 1, the $L_{1}$ inference incorporates one order of magnitude more non-zero edges ($E=10,911$) when compared with the MDL method ($E=1,540$). The edges obtained with $L_{1}$ are mostly of low magnitude (see Fig. 3c), and based on our analysis for synthetic data, are much more likely to be overfitting. We can find additional evidence of this by performing cross-validation also with the MDL method, as shown in Fig. 3d, which yields a significantly improved predictive performance of the held-out data. Both the predictive performance and description length improve when performing MDL regularization together with community detection, in accordance with Ref. [68].
@@
The interpretation of the two inferred models is fairly different. Although they share similar large-scale structures (as shown by the node colors indicating the inferred communities), the one inferred by $L_{1}$ possesses weights that are often negative, indicating some amount of direct antagonism between deputies, i.e. conditioning on the “yes” vote of one deputy will increase the probability of a “no” vote of another. Instead, the model inferred by MDL possesses only positive weights, divided in three categories, indicating an overwhelming tendency towards selective agreement, instead of direct antagonism. The groups it finds also align very well with the available metadata: the largest group corresponds to the governing coalition, whereas the remaining are opposition parties forming fragmented blocks  [^5] Both methods also lead to inferences of isolated nodes. These correspond to deputies who are mostly absent from voting sessions, and more typically abstain..
@@
Figure 4: Decimation procedure [35] employed on the same data as Fig. 3. The bottom panel shows the maximum likelihood growing monotonically as a function of the number of nonzero edges considered during decimation, and the top shows the similarity (weighted and binarized) with the network inferred via MDL. The vertical dashed line corresponds to the value $E=24,196$, obtained using the stopping criterion proposed in Ref. [35].
@@
We take the opportunity to compare with another regularization scheme called “decimation,” proposed in Ref. [35]. That approach is not based on shrinkage, and instead proceeds by first performing an unregularized maximum likelihood, resulting in a full network with no zero weights, then forcing a small number of the weights with smallest magnitude to zero, and proceeding recursively on the remaining entries in the same manner, until sufficiently many entries have been set to zero. Besides the increased computational cost of inferring the weights multiple times, and starting with a full network—thus imposing an overall complexity of at least $O(N^{2})$—this approach offers no obvious, principled criterion to determine when to stop the decimation. In Ref. [35] a heuristic stopping criterion was proposed by observing a point at which the likelihood no longer changes significantly in synthetic examples sampled from the true model. However, there is no guarantee that such a signal will exist in empirical data. In Fig. 4 we show for our Brazilian congress example the model likelihood offers no clear indication of when to stop the decimation and how many edges should be inferred in the end. The Jaccard similarity with the network inferred via MDL remains low through the whole procedure, peaking at close to $1/2$ for the binarized version when both networks have similar number of edges. When employing the heuristic stopping criterion proposed in Ref. [35], we obtain a network with $E=24,196$ edges — far denser that what we obtain with MDL. Due to these limitations, we see no advantage of using decimation over our approach based on MDL.
@@ VI Empirical case study: network of microbial interactions
Figure 5: Reconstructed networks for the (a) human microbiome project (HMP) with $N=45,383$, $M=4,788$, and $E=122,804$ and (b) earth microbiome project (EMP) with $N=126,730$, $M=23,323$, and $E=735,868$, using the MDL regularization method, together with a SBM prior. The top panels show the networks obtained, with edge weights indicated as colors. The middle panel shows the OTU counts in (a) different body sites and (b) earth habitats (the color code is the same as the “total count” panels in Fig. 9.). The bottom panel shows the edge weight distributions (with the same colors as the top panel), the node field distributions, the degree distributions (where node $i$ has total degree $k_{i}=\sum_{j}A_{ij}$, as well as positive and negative degrees, $k^{+}=\sum_{j}A_{ij}\mathds{1}_{W_{ij}>0}$ and $k^{-}=\sum_{j}A_{ij}\mathds{1}_{W_{ij}<0}$, respectively), the node strength distributions (where node $i$ has total strength $d_{i}=\sum_{j}W_{ij}$, as well as positive and negative strenght, $d_{i}^{+}=\sum_{j}W_{ij}\mathds{1}_{W_{ij}>0}$ and $d_{i}^{-}=\sum_{j}W_{ij}\mathds{1}_{W_{ij}<0}$, respectively), and the average node fields as a function of the degrees (with the expected “magnetization” in the inset).
@@
using the MDL regularization for both $\bm{W}$ and $\bm{\theta}$ discussed previously, including also the SBM prior. (We will not include a comparison with $L_{1}$, since the cross-validation procedure it requires becomes computationally prohibitive for problems of this magnitude.)
@@ VI.0.1 Predicting stability and outcome of interventions
In Fig. 7 we show the distribution of $z(j,\alpha)$ for macrostates simulated with the Ising model inferred from the HMP data, using the belief-propagation (BP) method [80], initialized with random messages. Although most perturbations are inconsequential, we observe that sometimes hundreds of species can disappear (or even appear) after a single species is removed.
@@ VII Conclusion
Acknowledgements.
This work has been funded by the Vienna Science and Technology Fund (WWTF) and by the State of Lower Austria [Grant ID: 10.47379/ESS22032]. It is also in part the result of research conducted for Central European University, Private University. It was made possible by the CEU Open Access Fund.
@@ Appendix A Generative models
In the case of the zero-valued Ising model with $\bm{x}\in\{-1,0,1\}^{N}$, but following the same Eq. 42, the normalization of Eqs. 43 and 45 change from $2\cosh(\cdot)$ to $1+2\cosh(\cdot)$.
@@ Appendix B Species taxonomies
Figure 9: Taxonomic classification of species for the human microbiome project (HMP) and the earth microbiome project (EMP), overlaid on the inferred networks of Fig. 5, together with the total species count over all samples.
In Fig. 9 we show the taxonomic classification of the species for the reconstructed networks of Fig. 5. The taxonomy consists of the phylum, class, order, family, and genus of each OTU. Each taxonomic category is shown as node colors, as indicated in the legend. In both cases we observe that most groups inferred with the SBM incorporated in our regularization scheme are compatible with the taxonomic division, since they tend to contain one dominating taxonomic category, at all levels — although a single category is often split into many groups, indicating structures that are not captured solely by taxonomy.